# Mean estimation in the add-remove model of differential privacy

### 11 Dec 2023 v1

@@ 3 Improving the Laplace mechanism via linear transformation
    Apply a linear transformation of matrix $\begin{bmatrix}1&-\frac{1}{2}\\ 0&\frac{1}{2}\end{bmatrix}$ to the [sum, count] vector.
@@
    Add vector Laplace noise according to the sensitivity of the new vector. This is equivalent to drawing the smallest $\ell_{1}$ unit ball in the transformed space for the K-norm mechanism.
@@
    Reverse the transformation by applying the inverse matrix $\begin{bmatrix}1&1\\ 0&2\end{bmatrix}$ to the noisy vector. Compute the noisy mean and truncate it.
Algorithm 2, along with the designed unit balls, is illustrated on the right of Figure 1. Compared with the red shape, this orange shape (no longer a square in the original 2D space) more tightly encloses the convex hull that we’d like to cover, hence reducing the Laplace noise scale.
The best possible seems to be fitting the unit ball exactly to the convex hull, which is the parallelogram. This is possible, since any linear transformation on a square results in a parallelogram. We only need to find a matrix, such that after multiplied with the matrix, the two segments $[(0,1),(1,1)]$ and $[-1,-1],[0,-1]$ are two sides of the $\ell_{1}$ square ball. This in turn gives Algorithm 3, claiming the smallest unit ball (and hence the Laplace noise scale) possible.
Figure 1: Comparison of Laplace mechanism tightest unit ball given different linear transformations. The middle figure shows the unit balls in the original 2D sensitivity space, where red=Algorithm 1; orange =Algorithm 2; blue=Algorithm 3, which is the smallest convex shape possible. The left figure shows the transformed space using the shifting in Algorithm 2; The right shows the transformed space using Algorithm 3.
Of course, the optimality of K-norm mechanism with the smallest convex hull is not guaranteed for any target metric. In the following sections, we show Algorithm 3 is in fact optimal for the mean squared error of the mean; and we get to release the count in add-remove model for free.
4 Analysis of algorithms
We first state a technical result which we use in proving the upper bound.
@@ 4 Analysis of algorithms
We provide the proof of the above result in the appendix. We now state the upper bound on the mean squared error of Algorithm 2.
@@
The output of Algorithm 2 is $\varepsilon$-differentially private. For any dataset $D\in\mathcal{D}^{*}(\ell,u)$, the mean squared error of Algorithm 2 is
$\mathbb{E}[(\hat{\mu}(D)-\mu(D))^{2}]$ $\leq\left(\frac{2(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{8(\mu-m)^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right)$
@@
Proof.
$\mathbb{E}[C^{2}]=\mathbb{E}\left[\left(\frac{Z_{s}}{n}-\frac{(\mu-m)Z_{n}}{n}\right)^{2}\right]=\frac{2(u-l)^{2}}{n^{2}\varepsilon^{2}}+\frac{8(\mu-m)^{2}}{n^{2}\varepsilon^{2}}.$
Since $(x+y)^{2}\leq 2x^{2}+2y^{2}$,
$\mathbb{E}[F^{2}]\leq 8\frac{(u-\ell)^{2}}{n^{4}}\mathbb{E}\left[Z^{4}_{n}\right]+8\frac{1}{n^{4}}\mathbb{E}\left[Z^{2}_{a}Z^{2}_{b}\right]=o\left(\frac{(u-l)^{2}}{n^{2}\varepsilon^{2}}\right).$
Finally, by the tail bounds of the Laplace mechanism,
$M^{2}\text{Pr}(Z_{b}<-b/2)\leq(u-\ell)^{2}\text{Pr}(Z_{n}<-n/2)=o\left(\frac{(u-l)^{2}}{n^{2}\varepsilon^{2}}\right).$
Combining the above three equations together with Lemma 1 yields the lemma. ∎
We now prove the upper bound on the mean squared error of Algorithm 3.
@@
The output of Algorithm 3 is $\varepsilon$-differentially private. Furthermore, for any dataset $D\in\mathcal{D}^{*}(\ell,u)$, the mean squared error of Algorithm 3 is
$\mathbb{E}[(\hat{\mu}-\mu)^{2}]$ $\leq\left(\frac{(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{4(\mu-m)^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right)$
@@
Let $s=[s_{1},s_{2}]$. $s$ is a two-dimensional vector. Let $s^{\prime}$ and $s^{\prime\prime}$ are vectors corresponding to two neighboring datasets. Then, the $L_{1}$ sensitivity is bounded by
$\|s^{\prime}-s^{\prime\prime}\|_{1}\leq w.$
In lines $5$ and $6$ of the algorithm, we add $\text{Lap}(w/\varepsilon)$ noise to each coordinate. Hence $[\hat{s}_{1},\hat{s}_{2}]$ is an $\varepsilon$ differentially private vector and by the postprocessing lemma, the output $\hat{\mu}$ is $\varepsilon$ differentially private. Similar to the previous lemma, the proof heavily relies in Lemma 1. Let $n=|D|$. Observe that
$\mathbb{E}[(\hat{\mu}-\mu)^{2}]=(u-\ell)^{2}\mathbb{E}\left[\left(\frac{\hat{s}_{1}}{\hat{s}_{1}+\hat{s}_{2}}-\frac{s_{1}}{n}\right)^{2}\right].$
Let $a=s_{1}$, $b=n$, $Z_{a}=Z_{1}\sim\text{Lap}(1/\varepsilon)$, $Z_{b}=Z_{1}+Z_{2}$, and $M=1$. With these definitions, to apply Lemma 1, we need to bound $\mathbb{E}[C^{2}]$, $\mathbb{E}[F^{2}]$, and $\text{Pr}(Z_{b}<-b/2)$. We bound each of the terms below.
$\mathbb{E}[C^{2}]=\mathbb{E}\left[\left(\frac{Z_{1}}{n}-\frac{(\frac{s_{1}}{n})(Z_{1}+Z_{2})}{n}\right)^{2}\right]=\frac{1}{n^{2}\varepsilon^{2}}+\frac{4(\frac{s_{1}}{n}-\frac{1}{2}))^{2}}{n^{2}\varepsilon^{2}}.$
@@
$\mathbb{E}[F^{2}]\leq\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{4}]+\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{2}Z^{2}_{1}]=o\left(\frac{1}{n^{2}\varepsilon^{2}}\right).$
Finally,
$\text{Pr}(Z_{b}<-b/2)\leq\text{Pr}(Z_{1}+Z_{2}<-n/2)\leq\text{Pr}(Z_{1}<-n/4)+\text{Pr}(Z_{2}<-n/4)=o\left(\frac{1}{n^{2}\varepsilon^{2}}\right).$
Combining the above three equations together with Lemma 1 and observing the fact that $\mu=\ell+\frac{s_{1}}{n}$ yields the theorem. ∎
The above theorem implies the following corollary.
@@
$R_{ar}(\varepsilon,\ell,u)\leq\frac{2(u-l)^{2}}{n^{2}\varepsilon^{2}}\left(1+o(1)\right).$

### 19 Feb 2024 v2

@@ 3 High privacy regime
    Apply a linear transformation given by the matrix $\begin{bmatrix}1&-\frac{1}{2}\\ 0&\frac{1}{2}\end{bmatrix}$ to the sum and count vector $q(D)$.
@@
    Add vector Laplace noise to the transformed vector according to its (reduced) sensitivity ($a=1$).
@@
    Reverse the transformation by applying the inverse matrix $\begin{bmatrix}1&1\\ 0&2\end{bmatrix}$, then divide to estimate the mean and truncate as before.
In the original (untransformed) $q$-space, depicted in the middle plot of Figure 1, the region protected by the resulting noise distribution is a parallelogram, shown in orange. Compared with the red diamond of Algorithm 1, this region more tightly encloses the sensitivity space, reducing noise and increasing accuracy. However, it still does not perfectly enclose $S(q)$.
3.1 Optimizing the Laplace mechanism via linear transformation
The approach in Algorithm 3 is to transform $q$ so that the $\ell_{1}$ unit ball encloses it as tightly as possible. To do this, it maps the two segments of the sensitivity space $[(0,1),(1,1)]$ and $[(-1,-1),(0,-1)]$ onto two sides of the $\ell_{1}$ ball using the transformation shown in the left plot of Figure 1. The enclosing $\ell_{1}$ ball is depicted in blue.
More concretely, Algorithm 3 computes the the scaled sum (denoted by $s_{1}$) and the difference of count and scaled sum (denoted by $s_{2}$). It then privatizes both $s_{1}$ and $s_{2}$ and uses them to compute mean. Next, we prove that Algorithm 3 instantiated with Laplace noise is not just intuitively better but in fact optimal for minimizing the mean squared error of the mean when $\varepsilon$ is sufficiently small.
3.2 Analysis of Laplace noise algorithms
We first state a technical result which we use in proving upper bounds.
@@ 3.2 Analysis of Laplace noise algorithms
$+2\sqrt{\mathbb{E}\left[C^{2}\right]\mathbb{E}\left[F^{2}\right]}$
[…]
We provide the proof in Appendix B. We next state an upper bound on the mean squared error of Algorithm 2.
@@
The output of Algorithm 2 is $\varepsilon$-differentially private. For any dataset $D\in\mathcal{D}^{*}(\ell,u)$, the mean squared error of Algorithm 2 is upper bounded by
$\left(\frac{2(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{8(\mu-m)^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right)$
@@
We provide the proof in Appendix C. We finally prove the upper bound on the mean squared error of Algorithm 3 when the noise distribution is Laplace.
@@
Let $Z_{1}$ and $Z_{2}$ be sampled from independent Laplace distributions with parameter $\varepsilon$. Then the output of Algorithm 3 is $\varepsilon$-differentially private. Furthermore, for any dataset $D\in\mathcal{D}^{*}(\ell,u)$, the mean squared error of Algorithm 3 is upper bounded by
$\left(\frac{(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{4(\mu-m)^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right)$
@@
Let $s=[s_{1},s_{2}]$ be a two-dimensional vector. Let $s^{\prime}$ and $s^{\prime\prime}$ be vectors corresponding to two neighboring datasets. Then the $\ell_{1}$ sensitivity is bounded by $\|s^{\prime}-s^{\prime\prime}\|_{1}\leq 1$.
In lines $5$ and $6$ of the algorithm, we add $\text{Lap}(1/\varepsilon)$ noise to each coordinate. Hence $[\hat{s}_{1},\hat{s}_{2}]$ is an $\varepsilon$ differentially private vector and, by the post-processing lemma, the output $\hat{\mu}$ is $\varepsilon$ differentially private.
The proof of the error bound relies heavily on Lemma 1. Let $n=|D|$. Since clipping only reduces the error,
$\mathbb{E}[(\hat{\mu}-\mu)^{2}]\leq(u-\ell)^{2}\mathbb{E}\left[\left(\frac{\hat{s}_{1}}{\hat{s}_{1}+\hat{s}_{2}}-\frac{s_{1}}{n}\right)^{2}\right].$
Let $a=s_{1}$, $b=n$, $Z_{a}=Z_{1}\sim\text{Lap}(1/\varepsilon)$, $Z_{b}=Z_{1}+Z_{2}$, and $M=1$. To usefully apply Lemma 1, we need bounds for $\mathbb{E}[C^{2}]$, $\mathbb{E}[F^{2}]$, and $\text{Pr}(Z_{b}<-b/2)$. We bound each of these below.
$\mathbb{E}[C^{2}]$ $=\mathbb{E}\left[\left(\frac{Z_{1}}{n}-\frac{(\frac{s_{1}}{n})(Z_{1}+Z_{2})}{n}\right)^{2}\right]$
$=\frac{1}{n^{2}}\mathbb{E}\left[\left(\left((1-\frac{s_{1}}{n}\right)Z_{1}-\frac{s_{1}}{n}Z_{2}\right)^{2}\right]$
$=\frac{1}{n^{2}\varepsilon^{2}}+\frac{4(\frac{s_{1}}{n}-\frac{1}{2}))^{2}}{n^{2}\varepsilon^{2}}.$ (6)
@@
$\mathbb{E}[F^{2}]$ $\leq\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{4}]+\frac{8}{n^{4}}\mathbb{E}[(Z_{1}+Z_{2})^{2}Z^{2}_{1}]$
$=o\left(\frac{1}{n^{2}\varepsilon^{2}}\right),$
where the last equality follows from the moments of the Laplace distribution (Kotz et al., 2012). Finally,
$\text{Pr}(Z_{b}<-b/2)$ $\leq\text{Pr}(Z_{1}+Z_{2}<-n/2)$
$\leq\text{Pr}(Z_{1}<-n/4)+\text{Pr}(Z_{2}<-n/4)$
$=o\left(\frac{1}{n^{2}\varepsilon^{2}}\right),$
where the last equality follows from the tail bounds of the Laplace distribution. Combining the three bounds above with Lemma 1 and observing the fact that $\mu=\ell+\frac{s_{1}w}{n}$ yields the theorem. ∎
Theorem 1 implies the following corollary, which shows that, in the high-privacy regime where $\varepsilon$ is small, the error of the add-remove model matches the swap model (Equation 3).
@@
[…]
4 Generalization to all values of $\varepsilon$
In this section we design an optimal $\varepsilon$-DP mechanism for private mean estimation in the add-remove model for any $\varepsilon$, dropping the high-privacy assumption. We prove the optimality of the new mechanism with respect to $R_{\text{ar}}(\varepsilon,\ell,u)$, and show that the optimal min-max error for the add-remove model is equivalent to that for the swap model, for any $\varepsilon$, up to a $(1+o(1))$ constant factor.
We first motivate the new mechanism, which we call the hourglass  mechanism. Observe that, in Algorithm 3, $s_{1}+s_{2}$ always sums to an integer, and hence the sensitivity space of $(s_{1},s_{2})$ is just the two bold segments on the left graph in Figure 1. Previously, we used the Laplace mechanism to protect the convex hull of these segments, but this is (still) a strict superset of the actual sensitivity space, which is non-convex. Figure 1(a) shows how many points in the $\ell_{1}$ unit ball actually cannot be reached by taking a single step to a neighboring database. The hourglass mechanism is designed to protect this sensitivity space more precisely, allowing for less noise. We formalize this notion and show that when the hourglass mechanism is used in Algorithm 3 to noise $[s_{1},s_{2}]$, the result is an optimal private mean estimator for all values of $\varepsilon$.
(a) The “hourglass” sensitivity space. Blue solid arrows show how, when moving to a neighboring database, the origin can shift only to points along the line segments $(0,1)$ to $(1,0)$ and $(0,-1)$ to $(-1,0)$. Red dashed arrows show how points like $(0.5,-0.5)$ are only reachable via a chain of two neighboring steps, even though though they lie within the $\ell_{1}$ unit ball.
(b) Illustration of the hourglass distribution for $\Delta=1$ and $\gamma=0.4$ in the range $[-4,4]^{2}$. All points on a line segment marked $j$ have density proportional to $e^{-j\varepsilon}$.
In particular, it is known that for privately computing one-dimensional sums (such as $s_{1}$ or $s_{2}$) the optimal mechanism is the staircase mechanism Geng and Viswanath (2014). The hourglass mechanism is constructed so that its marginal distributions both exactly match the univariate hourglass mechanism, with no change in $\varepsilon$.
4.1 Hourglass mechanism
The hourglass mechanism adds two-dimensional noise drawn from a distribution parameterized by $\gamma\in(0,1]$. Its density $f_{\gamma}(x,y)$ is supported on $\{(x,y):x+y=k\Delta,k\in\mathbb{Z}\}$. For $x\geq 0$, we divide each diagonal line $x+y=k\Delta$ into regions according to the value of $x$. For integers $i\geq 1$:
$A_{k}$ $:x\in[0,(k+\gamma)\Delta)$
$B_{k}(i)$ $:x\in[(k+\gamma+i-1)\Delta,(k+\gamma+i)\Delta)$
Note that, for $k<0$, $A_{k}$ is always empty, and $B_{k}(i)$ does not contain any points with $x\geq 0$ unless $i\geq-k$.
The density of the hourglass noise distribution is given by
$f_{\gamma}(x,y)\propto\begin{cases}e^{-k\cdot\varepsilon},&(x,y)\in A_{k}\\ e^{-(2i+k)\cdot\varepsilon},&(x,y)\in B_{k}(i)~{}.\end{cases}$ (7)
For $x<0$, $f_{\gamma}(x,y)=f_{\gamma}(-x,-y)$. See Figure 1(b) for an illustration of the hourglass distribution when $\Delta=1$ and $\gamma=0.4$.
Note that the hourglass mechanism is different from the natural extension of the univariate staircase mechanism to two dimensions as proposed by Geng et al. (2015). The latter is known to be optimal for an $\ell_{1}$-ball sensitivity space, but does not have the properties shown in Lemma 3, which are key to the optimality of the hourglass mechanism, and does not perform as well in practice (see Section 6).
Theorem 2.
Let $q:\mathcal{D}\to R^{2}$ be a query such that for any two neighboring datasets $D$ and $D^{\prime}$,
$q(D)-q(D^{\prime})\in\{(x_{0},\Delta-x_{0}):x_{0}\in[0,\Delta]\}.$
Then the hourglass mechanism given by
$q(D)+(Z_{1},Z_{2}),$
where $Z_{1},Z_{2}$ are sampled according to the density $f_{\gamma}$, is $\varepsilon$-DP.
Proof.
We provide an intuition based on Figure 1(b) here and a detailed proof in Appendix D.2. Recall that a segment marked by integer $j$ in Figure 1(b) has density proportional to $e^{-j\varepsilon}$, and on neighboring databases the noise distribution is effectively shifted by $(x_{0},1-x_{0})$ for some $x_{0}\in[0,1]$. Thus, to ensure privacy, it must be the case that the integers marking any pair of points $(x,y)$ and $(x+x_{0},y+1-x_{0})$, which are on adjacent lines in Figure 1(b) and not more than one unit apart along either axis, differ by at most one. By construction, this is true across the support of $f_{\gamma}$. ∎
We now show that the marginal distributions of the hourglass mechanism are univariate staircase mechanisms, which is the key ingredient to proving optimal utility guarantees.
Lemma 3 (Marginal distribution).
The marginal densities $f_{\gamma}(x)$ and $f_{\gamma}(y)$ match the $\varepsilon$-DP univariate staircase mechanism with parameter $\gamma$. That is, $f_{\gamma}(x)\propto e^{-\varepsilon\lceil|x|-\gamma\rceil}$, and $f_{\gamma}(y)\propto e^{-\varepsilon\lceil|y|-\gamma\rceil}$.
Proof.
Assume $x\geq 0$, and let $j\geq 0$ be the unique integer for which $x\in[j-1+\gamma,j+\gamma)$. Then the marginal density for $x$ is
$f_{\gamma}(x)$ $\propto\sum_{k=-\infty}^{\infty}f_{\gamma}(x,k-x)$ (8)
$=\sum_{k=-\infty}^{j-1}e^{-(2j-k)\varepsilon}+\sum_{k=j}^{\infty}e^{-k\varepsilon}$ (9)
The first term follows because $j-k\geq 1$ in this range, and so $x\in B_{k}(j-k)$. The second term follows because $x\in A_{k}$ in this range. Summing the geometric series, the above is equal to
$e^{-(j+1)\varepsilon}/(1-e^{-\varepsilon})+e^{-j\varepsilon}/(1-e^{-\varepsilon})\propto e^{-j\varepsilon}~{},$
which is precisely the staircase density. The densities for $x<0$ and $y$ are handled symmetrically. ∎
Next we derive the conditional distribution of $y$ given $x$.
Lemma 4 (Conditional distribution).
The conditional probability under $f_{\gamma}$ of $y$ given a fixed $x$ is a geometric distribution with ratio $e^{-\varepsilon}$. In particular, for any $y$ such that $y+x$ is an integer,
$f_{\gamma}(Y=y|x)=\frac{1-e^{-\varepsilon}}{1+e^{-\varepsilon}}e^{-\varepsilon|y-y_{0}(x)|},$
where
$y_{0}(x)=\begin{cases}-x+\lfloor x+(1-\gamma)\rfloor,&x\geq 0\\ -x-\lfloor-x+(1-\gamma)\rfloor,&x<0\end{cases}$ (10)
Using the above lemmas, we get a simple sampling algorithm for the hourglass distribution: first sample $x$ from a univariate staircase mechanism with $\gamma$ and $\varepsilon$, and then sample $y$ from the geometric distribution in Lemma 4.
4.2 Implications for mean estimation
Before proceeding to analyze the accuracy of the hourglass mechanism for mean estimation, we first consider the two-dimensional staircase mechanism.
Lemma 5.
Let $Z_{1}$ and $Z_{2}$ be sampled by the two-dimensional staircase mechanism with parameter $\varepsilon$ and with parameter $\gamma$ optimized for mean squared error. Then the output of Algorithm 3 is $\varepsilon$-DP. Furthermore, for any dataset $D\in\mathcal{D}^{*}(\ell,u)$, the MSE of Algorithm 3 is upper bounded by
$\left(\frac{(u-l)^{2}\widetilde{\sigma}^{2}(\varepsilon)}{|D|^{2}}\right)\left(1+o(1)\right),$
where $\widetilde{\sigma}^{2}(\varepsilon)$ is the variance of the two-dimensional staircase mechanism optimized for MSE.
The proof is provided in Appendix E. We now state our main upper bound for all values of $\varepsilon$. The proof is similar to that of Lemma 5 together with the fact that the marginal distribution along each dimension is the same as the univariate staircase mechanism (Lemma 3). We provide the full proof in Appendix F.
Theorem 3.
Let $Z_{1}$ and $Z_{2}$ be sampled by the hourglass mechanism with parameter $\varepsilon$ and $\gamma$ as given by Geng and Viswanath (2014, Equation 50). Then the output of Algorithm 3 is $\varepsilon$-differentially private. Furthermore, for any dataset $D\in\mathcal{D}^{*}(\ell,u)$, the mean squared error of Algorithm 3 is upper bounded by
$\left(\frac{(u-l)^{2}{\sigma}^{2}(\varepsilon)}{|D|^{2}}\right)\left(1+o(1)\right),$
where $\sigma^{2}(\varepsilon)$ is defined in (2).
(c) All Laplace mechanisms, high privacy regime, varying $\varepsilon$.
(d) All Laplace mechanisms, high privacy regime, varying $\mu$.
(e) Hourglass and 2D staircase mechanisms, low privacy regime, varying $\varepsilon$.
Figure 2: Error comparison of different algorithms with varying $\varepsilon$ or $\mu$.
@@ 5 Lower bound
We next state a well-known result of Geng and Viswanath (2014) that we will use to prove our lower bound.
Lemma 6 (Geng and Viswanath (2014, Section VI.C)).
Let $k>0$ and let $\mathcal{D}^{k}$ denote the collection of all datasets of size $k$ where each value lies in the range $[0,1]$. For $D\in\mathcal{D}^{k}$, let $S(D)$ denote the sum of the elements in $D$. Let $\hat{S}$ denote an estimator of the same sum. Then, for any $\varepsilon$,
[…]
where $\sigma^{2}(\varepsilon)$ is given by (2). Here the $o(1)$ term goes to zero as $k$ tends to infinity.
Lemma 6 provides a lower bound for estimating the sum over datasets of fixed size. We construct a class of datasets $\mathcal{D}^{\prime}$ with varying size such that for every dataset in $\mathcal{D}^{\prime}$ there exists a corresponding dataset in $\mathcal{D}^{k}$ (as defined in Lemma 6) with the same sum. Hence, any differentially private mean estimator for the datasets in $\mathcal{D}^{\prime}$ can be modified to obtain a differentially private sum estimator for the datasets in $\mathcal{D}^{k}$. We use this observation to show the following lower bound on the min-max MSE of mean estimation under the add-remove model of differential privacy.
Theorem 4.
[…]
Combining Theorem 3 together with Theorem 4 and (2) yields the following result.
Corollary 2.
[…]
$\frac{2^{-2/3}e^{-2\varepsilon/3}(1+e^{-\varepsilon})^{2/3}+e^{-\varepsilon}}{(1-e^{-\varepsilon})^{2}}\left(1\pm o(1)\right).$
Furthermore, the hourglass mechanism with suitable value of $\gamma$ achieves this normalized MSE.
6 Experiments
In Figure 2 we plot the empirical performance of the algorithms discussed in Sections 3 and 4.1 on synthetic datasets and explore how the performance changes with parameters such as the privacy budget $\varepsilon$ and the true mean $\mu$. The underlying datasets are generated i.i.d. with varying $\mu$ in the range $[\ell=0,u=1]$. All datasets have $10,000$ points, and mean squared error is computed over $100,000$ runs of each algorithm. The errors are normalized by $\frac{|D|^{2}\varepsilon^{2}}{2}$ to keep them in a similar range across a wide range of parameters.
Figures 1(c) and 1(d) compare the performance of Algorithms 1, 2 and 3 using the Laplace mechanism, varying $\varepsilon$ and $\mu$, respectively. These plots focus on the high-privacy regime, where Algorithm 3 was shown to be optimal. Indeed, Algorithm 3 outperforms the others, reducing error over Algorithms 2 by roughly a factor of two, which matches the ratio of upper bounds in Lemma 2 and Theorem 1. Similarly, in Figure 1(c), Algorithm 3 closely matches the lower bound $\sigma^{2}(\varepsilon)$ from (2) when $\varepsilon<1$.
In Figure 1(d), we explore how the error changes with the true mean of the data. As analyzed in (3.2), the error is largest when $\mu$ approaches 00 or $1$. However, the error also drops when $\mu$ becomes very close to 00 or $1$, since $\frac{\hat{s_{1}}}{\hat{n}}$ falls out of the range $[\ell,u]$ with probability approaching 50%. In these cases the clipping operation significantly reduces the mean squared error.
Finally, we compare the two-dimensional staircase mechanism to Algorithm 3 with noise from the hourglass mechanism in Figure 1(e), focusing on the low-privacy regime where $\varepsilon$ is large. The lower bound in Figure 1(e) is again $\sigma^{2}(\varepsilon)$, but notice now that as $\varepsilon$ grows, the hourglass mechanism continues to match the lower-bound indefinitely, and increasingly outperforms the two-dimensional staircase mechanism. In Figure 3 (Appendix H), we show that both mechanisms also have a similar M-shaped error curve over $\mu$. In Appendix H, we provide additional experiments to demonstrate that the proposed algorithm outperforms existing algorithms on small datasets.
7 Acknowledgements
The authors thank Travis Dick, Gautam Kamath, and Ziteng Sun for helpful comments and suggestions.
Appendix A Properties of staircase mechanisms
We first define both the one-dimensional and two-dimensional staircase mechanisms.
Definition 2 (Univariate staircase mechanism).
Let $\Delta$ be the sensitivity of the underlying query. The univariate staircase mechanism Geng and Viswanath [2014] is parameterized by $\gamma\in[0,1]$ and is given by
$f_{\gamma}(x)=\begin{cases}a(\gamma),&x\in[0,\gamma\Delta)\\ a(\gamma)e^{-\varepsilon},&x\in[\gamma\Delta,\Delta)\\ e^{-k\varepsilon}f_{\gamma}(x-k\Delta),&x\in[k\Delta,(k+1)\Delta)\\ f_{\gamma}(-x),&x<0.\end{cases}$ (11)
Here $a(\gamma)$ is the normalization factor and given by
$a(\gamma)=\frac{1-e^{-\varepsilon}}{2\Delta\left(\gamma+e^{-\varepsilon}(1-\gamma)\right)}.$
Definition 3 (Two-dimensional staircase mechanism).
Let $\Delta$ be the sensitivity of the underlying query. The two-dimensional staircase mechanism Geng et al. [2015] is parameterized by $\gamma\in[0,1]$ and is given by
$f_{\gamma}(x,y)=\begin{cases}a(\gamma)e^{-\varepsilon k},&|x|+|y|\in[k\Delta,(k+\gamma)\Delta)\\ a(\gamma)e^{-\varepsilon(k+1)},&|x|+|y|\in[(k+\gamma)\Delta+k+1)\Delta),\end{cases}$ (12)
Here $a(\gamma)$ is the normalization factor and given by
$a(\gamma)=\frac{(1-e^{-\varepsilon})^{2}}{2\left(e^{-\varepsilon}(e^{-\varepsilon}+(1-e^{-\varepsilon})\gamma)+(1-e^{-\varepsilon})(e^{-\varepsilon}+(1-e^{-\varepsilon})\gamma^{2})\right)}.$
We prove the following result for the MSE of the two-dimensional staircase mechanism.